OpenStudy (anonymous):

Someone help. Let f(x,y,z) be a function of three variables. Suppose that C is an oriented curve lying on the level surface f(x,y,z)=1. Find the line integral ∫∇f⋅dr ⃗ and explain your answer.

OpenStudy (amistre64):

can you write the integral out using the draw box or the equation tool? hard to tell what [] is spose to mean

OpenStudy (anonymous):

Find the line integral \[\int\limits_{a}^{b} ∇f⋅dr ⃗ \]

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

that does look better :) how far have you gotten so far? or what about the process are you stuck at?

OpenStudy (anonymous):

i am kind of thinking of fundamental theorem of line integral but not sure about it

OpenStudy (amistre64):

my first thought was stokes, but i am a bit rusty

OpenStudy (anonymous):

whatever help would be good. What do you think? I mean i have no idea so anything would be helpful.

OpenStudy (amistre64):

my difficulty at the moment is try to interpret the information, it sounds very general. \[f(x,y,z)=1\] \[f_x=0~:~f_y=0~:~f_z=0\] \[\nabla f=(f_x,f_y,f_z)=(0,0,0)\] the dot product to this would just be 0, so I would say that the integration is up to a variable: t \[\int_{a}^{b} 0 ~dt=b-a\]

OpenStudy (amistre64):

but then again, i am a it rusty and i cannot verify this to be correct

OpenStudy (amistre64):

http://openstudy.com/study#/updates/52a528c7e4b0256dd0989f74 this person is asking the same question ... you may be able to help each other out.

OpenStudy (anonymous):

\[ f(x1,y1,z1) - f(xo, yo,zo) = 1-1= 0\] that what i was thinking of .... using fundamental theorem.

OpenStudy (anonymous):

@amistre64 yeah we have exact same question.... hopefully i might get some help from him

OpenStudy (amistre64):

good luck with it :) if i come across anything that I can be sure about, I will let you know.

OpenStudy (anonymous):

@amistre64 thanks...